OpenStudy (anonymous):

I need help please, Let f(x) = x – 3 and g(x) = x + 11. Find f(x) • g(x). x2 – 8x – 33 x2 – 8x + 14 x2 + 8x + 14 x2 + 8x – 33

OpenStudy (anonymous):

Just multiply (x-3)(x+11) You will have x² - 3x + 11x - 33 = x² + 8x - 33

OpenStudy (anonymous):

do you do alike terms?

OpenStudy (anonymous):

I really need help anyone.

OpenStudy (jdoe0001):

just straightforward multiplication, element by element

OpenStudy (jhannybean):

\[f(x)=x-3\]\[g(x)=x+11\]\[f(x) \cdot g(x) = (x-3)(x+11)\]\[\color{green}x(x+11) -\color{purple}3(x+11)\]\[x(x)+11(x) -3(x)-3(11)\]Can you simplify this now?

OpenStudy (jdoe0001):

(a+b) (c-d) a(c-d) + b(c-d)

OpenStudy (anonymous):

(x2+11x)(-3x+8) so far correct?

OpenStudy (anonymous):

wouldnt it be the 3rd one?

OpenStudy (jhannybean):

No, what is -3 x 11?

OpenStudy (anonymous):

-33

OpenStudy (anonymous):

so the last one